OpenStudy (hannahduhe):

Given the information below, find the coordinates of the vertices L and P such that ABCD IS similar to NLPM A(2,0), B(2,4), C(-2,4) D(-2,0) M(4,0) N(12,0) Answers A)L(8,4), P (8,12) B)L(12,8), P(4,8) C)L(4,6), P(12,6) D)L(6,8), P(8,8)

OpenStudy (trojanpoem):



OpenStudy (trojanpoem):

AD/AB = NM/NL 4/4 = 8/y 1 = 8/y y = 8 so P = (4,8) , L = (12, 8)